Qd = 1500 - 5P

Qs = -500 + 5P

Find

P, Q


1
Expert's answer
2021-02-09T07:01:52-0500

At the equilibrium "Q_d=Q_s", therefore, we can write:


"1500-5P=-500+5P,""2000=10P,""P=200."

Then, we can substitute "P" into the demand function and find "Q":


"Q=1500-5\cdot200=500."